A traffic policeman standing on a road sounds a whistle emitting the main frequency of 2.00 kHz. What could be the apparent frequency heard by a scooter-driver approaching the policeman at a speed of 36.0 km h-1? Speed of sound in air = 340 m s-1.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Here given, apparent frequency = 1620 Hz So original frequency of the train is given by 1620 = (332 + 0 + 0/332 - 15) f ⇒ f = (1620 * 317/332) Hz So, apparent frequency of the train observed by the observer in F^1 = (332 + 0 + 0/332 + 15) f * (1620 * 317/332) = 317/347 * 1620 = 1480 Hz.